OpenStudy (anonymous):

Linear Algebra: Determine whether the matrices At + A and At - A are symmetric or skew-symmetric.

OpenStudy (anonymous):

\[A^{T} - A, A^{T} + A\] "T" referring to the transpose of A

OpenStudy (helder_edwin):

what do u get when u do the following? \[ \large (A^t-A)^t= \] and \[ \large (A^t+A)^t= \]

OpenStudy (helder_edwin):

remember this \[ \large (A+B)^t=A^t+B^t \] \[ \large (rA)^t=r\cdot A^t \]

OpenStudy (anonymous):

\[(A^{T} - A)^{T} = A - A^{T}, (A^{T} + A)^{T} = A + A^{T}, \]

OpenStudy (helder_edwin):

great. now u know this definitions: \[ \large A\text{ is symmetric if }A^t=A \] \[ \large A\text{ is skew-symmetric if }A^t=-A \]

OpenStudy (anonymous):

then I just add or subtract to both sides?

OpenStudy (helder_edwin):

no

OpenStudy (helder_edwin):

u got this \[ \large (A^t+A)^t=(A^t)^t+A^t=A+A^t=A^t+A \] so according with the definition i gave u. is this matrix symmetric or skew-symmetric?

OpenStudy (helder_edwin):

well @jcd2012 ?

OpenStudy (anonymous):

im trying to formulate an answer, you're basically saying I should think as the transpose of the sum and differences as one matrix being transposed, so it meets the definition?

OpenStudy (helder_edwin):

exactly

OpenStudy (anonymous):

then by the definition, (A^T - A) is skew-symmetric and (A^T + A) is symmetric

OpenStudy (anonymous):

right?

OpenStudy (helder_edwin):

right.